OpenStudy (curry):

Last probability question for the entire quarter (for hw atleast, still have a final to go, yay!)

ganeshie8 (ganeshie8):

Hey still here ?

ganeshie8 (ganeshie8):

@Curry

OpenStudy (curry):

yes

OpenStudy (curry):

@ganeshie8

OpenStudy (curry):

Hey ganesh, quick question! P(A | B) = P(A) can only be true if P(B) is null right? So that when P(A|B) = P(A) | P(B) = P(A)?

OpenStudy (curry):

here's what i have for this question so far, F = spam, E = herbal, P(F|E) = P(E|F) P(F)

ganeshie8 (ganeshie8):

that algebra confuses me

ganeshie8 (ganeshie8):

lets draw a tree diagram

ganeshie8 (ganeshie8):

are you good with tree diagrams ?

OpenStudy (curry):

Um, yes i am!

ganeshie8 (ganeshie8):